Tech Staff Meeting Notes January 18, 2006

Current Top Priorities are:Station Maintenance, Broadband Upgrade

Steve, Jim, and Mitch will be participating in the ANSS NetOps workshop February 6-7 in Salt Lake City.

Feast of Flesh is February 16 at 6pm at the Butcher Shop, Cordova.


There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about


The efforts to fill the open Research Equipment Tech position continue. This redistributed task list agreed to at the November meeting. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.

Greg's October ToDo List

Priorities

BB network upgrade.
Intuicom mockup is working. Next step is to try it in the field at a realistic distance. Breakouts are built and Diag board is nearing completion (for temp, alarm, etc). Steve has tank data and the McNamara PDF software running. Need to convert to mseed and get generic dataless seed for CMG3/DM24 and CMG40T/DM24.

GPS
Columbia, TN permission is a wash. Still don't have an eastern hard rock station.
Need to notify James when we do the fiber isolation at the CVTN strongmo station so he can verify we didn't muck up the GPS station.
Camp Robinson needs a routine service. Contact info needs to be updated and is likely best done on site.
Need to test aerocom telemetry path at Steele. Station is at private business near airport and the plan is to telem to High School on the other side of town where we can get internet.
Piggot not yet converted to ethernet though permission is secured. IT tech is difficulty to contact.
Need a new plan for MS County Airport at East Prairie, MO.
Reelfoot needs to be converted from Solar to AC.

ETN

John sent an updated planned, in progress, and completed station list and map. Mitch will add strongmo and bb. Then need to meet to develop long term plan. This is needed to complete schematics required for the Mt Mitchell node agreement.
Gate at Van Hill complete.
John will be visiting Memphis in the next week to month. He will stop by LRNC and bring the rack back. He and Greg will work together to rebuild it in preparation for installation at Mt. Mitchell.

ANSS.
Jim took over the ANSS stations. Is working on getting all the records straightened out.
HAIL and SEAR still need to be swapped with the 4g (+-2g).
Continued problems with mss100's losing connectivity (SMOC: simple matter of connectivity).
Mitch has the program to fix the leap second error. Need to figure out which ones actually need fixing. Some stations will require a site visit for visual verification. Holly will look for events from 8/4-12/31 to see if we can verify using data.
Need to upgrade trimble firmware to avoid leap second error next time. A workable upgrade procedure doesn't yet exist. Jim, Wurth, and DTA or working on it.
CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. Filipcic needs to install. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small).

Routine Maintenance. Jim knows what needs work.

Power for V890 nearly complete. Need to pay greater attention to Liebert warning since hardware we now have will require both compressors during hot weather. Dan at P-plant once had plans to monitor remotely. Should we do this (for ups's too?)?

Emergency lights in grad building don't work except for the ones in the hallway.


Second Tier Priorities

I-40 Bridge.
W28 borehole sensor has a bad component.
A large piece of concrete fell and crushed the conduit to one of the sensors. Will need to pull new wires and install new flex conduit.
All DAS's were operational for 22 continuous days and we have the data. Four are now not working. Steve thinks the disks filled up and is working on a batch file to clean up old files. Also, it looks like some don't get gps lock before the DAS times out and reboots. Greg has a fix for this.

FDL. We'll start keeping fdl notes in a separate fdl section.
Don't see a need to do that now that Jim took over and changed my mind.
Made excellent progress in cleaning up. Still need to get more "tubs" for tools and small items.
Where are all the laptops?
Checkout/in procedures/forms?
Need to clean brushes and check connectors on resistivity meter.


As Time Permits


Other

NHIN.
Needs a new dsl modem.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.

Timing. Steiner has an easy way to generate a datastream that goes high when lock is lost. Need a mask to decide what streams to use. Need mechanism to produce alarms. BB stations more difficult. How do we automate the status stream or even get it back to Memphis for logging? (Save it to file with Scream?). Note, new version of firmware sends status to both flash and stream. Hence, information will be in gcf2ew logs, but how do we monitor that? Maybe not a huge deal with broadbands since both Bollwerk and Brewer regularly monitor. Would still be good to have a record though.

VBMS. Everything is in order and ready to go. Installation will be this fall depending on schedule of contractor. Greg may visit during installation if schedule permits.

Langston student projects continue. See Greg's list above.

Shop reorganization nearly complete.